OpenStudy (anonymous):

How much of 5/6 is in 1/3?

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (anonymous):

2/5 I think?

OpenStudy (anonymous):

But how do you get it?

OpenStudy (dmndlife24):

WN x 5/6 = 1/3 WN5/6 = 1/3 WN = 2/5 Yes you are correct!

OpenStudy (anonymous):

Thank You!